OpenStudy (anonymous):

Find the center and radius of the circle whose equation is x^2 -8 x + y^2 + 6 y -9 = 0

OpenStudy (anonymous):

The equation for a circle is (x-h)^2+(y-k)^2=r^2, where the center is (h,k)

OpenStudy (anonymous):

Simplify it to (x-h)^2+(y-k)^2=r^2